OpenStudy (anonymous):

Simplify the expression. Write the answer without using parentheses or negative exponents. x^-3

OpenStudy (anonymous):

@professional_loser_

OpenStudy (anonymous):

@dpasingh

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \sf a^{-b} =}\)\(\Huge\color{blue}{ \sf \frac{1}{a^b} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \sf x^{-3} =}\)\(\Huge\color{blue}{ \sf \frac{?}{?} }\)

OpenStudy (anonymous):

1/x^3 is the answer. just like the last problem, take the portion with the negative exponent to the bottom, and you have your answer of 1/x^3.

OpenStudy (anonymous):

WILL GIVE MEDAL

OpenStudy (anonymous):

\[\huge x^{-3} = \frac{1}{x^3}\]

OpenStudy (anonymous):

@cancherolugano

OpenStudy (anonymous):

??????